What is the strain energy stored in a body due to bending moment? *


1
Expert's answer
2021-01-13T11:37:16-0500

Let's consider a small segment of beam of length "ds" subjected to bending moment. One cross section rotates about another cross section by a small amount "d\theta" and we can write:


"d\theta=\dfrac{1}{R}ds=\dfrac{M}{EI}ds, (1)"

here, "R" is the radius of curvature of the bent beam, "M" is the bending moment, "EI" is the flexural rigidity of the beam.

Let's write the work done by the bending moment while the segment of beam rotating through angle "d\theta":


"dU=\dfrac{1}{2}Md\theta (2)"

Substituting formula (1) into the formula (2), we get:


"dU=\dfrac{1}{2}\dfrac{M^2}{EI}ds (3)"

Finally, we can find the strain energy stored in the beam of span "L" by integrating equation (3) over "L":


"U=\displaystyle\intop_{0}^L \dfrac{M^2}{2EI}ds."

Answer:

"U=\displaystyle\intop_{0}^L \dfrac{M^2}{2EI}ds."